OpenStudy (anonymous):

A person walks first at constant speed of 5.00 m/s along a straight line from point A to point B and then back along the line from B to A at a constant speed of 3.00 m/s. What is her average speed over the entire trip?

OpenStudy (ajprincess):

Let the distance between A and B be d. Speed s=d/t 5=d/t1 t1=d/5 3=d/t2 t2=d/3 Average speed=(d+d)/(t1+t2) Hope that helps:)

OpenStudy (anonymous):

aveg. speed = total distance / total time let x be the distance from A to B. The distance from B to A is also x. So the total distance is 2x time = distance / speed time from A to B: t = x/5 time from B to A t = x/3 the total time is x/5 + x/3 = 8x/15 averg. speed = 2x / (8x/15) averg. speed = 3.75 m/s aveg. velocity = total displacement / total time to find displacement, take final postion and subtract the initial position the inital position is 0m. the finial position is also 0m, because the person goes back to here he startedd total displacement is 0 - 0 = 0m aveg. velocity = 0 / (8x/15) = 0m/s hope this helps